# Towards an Accessible, Noninvasive Micronutrient Status Assessment Method: A Comprehensive Review of Existing Techniques

### v1

@@ 2.1 Micronutrient Characteristics
Micronutrients are divided into three categories: water-soluble vitamins, fat-soluble vitamins, and minerals [24]. Water-soluble vitamins, such as vitamin C and the B vitamins, are absorbed directly into the bloodstream and are quickly excreted in urine. They are not stored for long periods in the body, so regular intake is necessary to prevent deficiencies, and there is less concern about toxicity from excess intake. On the other hand, fat-soluble vitamins like A, D, E, and K are absorbed into lymph vessels along with dietary fats. Fat-soluble vitamins are stored in larger quantities in fatty tissues and the liver, so deficiencies take longer to develop, and daily intake is less critical. However, due to their efficient storage and the lack of a rapid excretion mechanism, toxicity is more of a concern.
@@
Minerals can be categorized as major or trace minerals based on the daily requirement. Major minerals, such as sodium, potassium, chloride, phosphorus, and magnesium, are required in amounts greater than 100 mg per day, while trace minerals like iron, copper, zinc, selenium, iodine, chromium, fluoride, and manganese are needed in amounts of 100 mg or less per day. Minerals are water-soluble and are absorbed directly into the bloodstream, sometimes with the help of transport proteins. It is important to note that minerals have an electric charge, and their function and storage can be influenced by various factors. A comprehensive summary of the characteristics of micronutrients is presented in the Appendix in Tables LABEL:tab:char-vitamins-water-soluble, LABEL:tab:char-vitamins-fat-soluble, and LABEL:tab:char-minerals.
@@
Dietary guidelines for intake vary across organizations and countries and are based on factors such as age and sex. These guidelines are not standardized due to the individualized nature of micronutrient metabolism, the diversity of micronutrients, their interactions, and the specific values at which they are needed. In the US, dietary reference intakes are used, mostly utilizing the Recommended Dietary Allowance (RDA). The RDA represents the average daily level of intake sufficient to meet the nutrient requirements of 97-98% of healthy individuals. The National Institutes of Health also defines terms to describe intake levels, such as an Upper Limit (UL), highlighting the potential for excess intake [120]. In cases where there is insufficient evidence to establish these guidelines, there exists the more general term, Adequate Intake, further emphasizing the uncertainty surrounding the appropriate levels of micronutrients in diet.
Micronutrient status assessment usually relies on the combination of the analysis of micronutrient biomarkers in biofluids as well as the physiological symptoms presented by an imbalance. Micronutrient imbalances have well-documented physical symptoms, which provides opportunities to apply noninvasive methods to determine micronutrient status [68, 109, 45, 39, 140]. These symptoms provide noninvasive insights into internal processes and help identify appropriate biomarkers for further testing. Micronutrient biomarkers are still debated, but can generally serve as a reliable gold standard for assessment methods [43].
2.2 Micronutrients in Biofluids
@@
Biofluid analysis for micronutrient status remains challenging since it is often unclear how other biofluids reflect the gold-standard matrix for that micronutrient. Some biofluids include blood, saliva, sweat, tears, and urine. Besides blood, each of these can be collected and analyzed noninvasively. Blood and urine are clinically relevant for representing in-body micronutrient status, as will be demonstrated in Section 3.1. However, evidence is less clear for saliva, sweat, and tears [69].
In the clinical literature, the equivalence of saliva, sweat, and tears to blood and urine in micronutrient status assessment remains ambiguous. Some evidence for correlation of micronutrient levels in saliva with blood levels was found on a by-micronutrient basis [69]. For example, serum and saliva levels of vitamin D were found to have a correlation of 0.56 [11]. Additionally, some correlation of iron levels in saliva and serum (a blood derivative) have been found. However, validity remains inconclusive as some sources have found a high positive correlation between salivary and serum level [26, 60], while others report a high negative correlation [49, 10, 53]. Despite this, research still exists that aims to quantify micronutrient status with saliva as well as tears [145, 150]. Sempionatto et al. [150] in particular argue that tears are a good biofluid for analysis since they are noninvasive, less complex than blood yet still contain a "variety of biomarkers", and they "reflect concurrent blood levels" because of passive leakage of compounds from blood plasma. It is important to note, however, that neither of these works provide comparison to the in-body status of their target micronutrients as measured by gold-standard, urine or blood-based assays.
@@
The relevance of sweat for micronutrition seems to have been the subject of more research. Indeed, sweat receives significant attention in emerging micronutrient detection methods, especially in wearables. When we discuss sweat, we refer to eccrine sweat. Eccrine sweat is mostly sodium and chloride, with smaller amounts of micronutrients and metabolites (at the micro and nanomolar scale), similar to or smaller than their concentrations in blood plasma [14]. Micronutrients found in sweat include potassium, calcium, magnesium, iron, copper, zinc, vitamin C, and vitamin B1. Out of the compounds in sweat, mostly sodium and chlorine ions are well-studied. Some research has explored water-soluble vitamins in sweat, such as vitamin C and vitamin B1, but there is no such attention on fat-soluble vitamins. Many confounding factors can impact sweat composition, such as whether the sweating is passive or active. Active sweating usually produces higher concentrations of sodium, chloride, and metabolites. Contamination of sweat by skin-derived substances, notably iron, can affect composition as well. This is combated by pre-rinsed skin, removal of initial sweat (concentrations stabilize after 20 to 30 minutes of sweating), and the analysis of cell-free sweat. The region and method of collection can also have an impact, as many micronutrient concentrations can vary two to four times depending on the region. Finally, sweat can be reabsorbed into the body. Skin temperature and the flow rate of sweat both impact the rate of this reabsorption.
@@
Unfortunately, the clinical literature notes a general "lack of association between dietary micronutrient intake and corresponding sweat micronutrient concentrations" [14]. This lack of an association exists in comparison to blood as well. A review by Baker and Wolfe [14] finds that there is no established correlation between sweat and blood composition, and there is "little support for using sweat as a surrogate for blood". Concentrations of minerals in sweat are much more varied than in plasma, likely because minerals bind to carrier proteins in blood. The review also reports little to no correlation between sweat and blood concentrations of vitamin C and iron status. This finding for iron is echoed in another paper that found both iron and calcium have no correlation between sweat and blood concentrations [13]. However, it was reported in the same paper that iron concentrations in sweat have been observed to be lower in anemic patients and higher in patients undergoing iron therapy.
2.3 Micronutrient Effects on Physiological Processes
@@
Most of the physiological effects of micronutrients are related to the autonomic functions of the body. The literature documents the importance of micronutrients to the proper functioning of the nervous system [1, 54, 32]. For example, a study of vitamin B12 deficiency compared responses to a 60-degree passive head up tilt test between a control group, a vitamin B12 deficient group, and a group with diabetes mellitus [16]. They found that the deficient group had comparable autonomic neuropathy to the diabetic group. Our exploration of observable impacts of deficiencies on autonomic functioning found three main affected areas relating to biofluid composition and physiological effects: cardiac functioning, sleep, and general symptoms.
2.3.1 Cardiac Function
One main aspect of autonomic functioning affected by nutrient imbalance is cardiac functioning. Several studies explore the interaction between heart rate variability (HRV) and micronutrient deficiencies. Components of HRV are associated with parasympathetic (PNS) and sympathetic nervous system (SNS) activity [154]. High-frequency (HF) bands reflect PNS activity and correspond to the respiratory cycle, while low-frequency (LF) bands reflect PNS, SNS, and baroreceptor activity. Vitamin B12 deficiency is one of the most documented in terms of impact to HRV, with evidence that it lowers HRV overall, impacting sympathetic indices the most [8, 166, 102, 16]. Supplementation of B12 was also demonstrated to return HRV indices to a comparably normal state [8]. Deficiency of vitamin D was found to lower HRV as well [102]. Calcidiol (25(OH)D) levels, a form of vitamin D, was shown to be associated with the ratio of LF to HF HRV power [108]. This metric is sometimes called sympathovagal balance and is intended to be a measure of ’balance’ between SNS and PNS activity, but there has been debate over this interpretation [154]. Iron-deficiency anemia (IDA), an advanced form of iron deficiency, has more conflicting evidence of HRV impacts, with some studies finding no difference in HRV indices versus the control [171] while others were able to find a difference in the IDA group [75, 188].
Impacts of micronutrients on blood pressure have also been studied [32, 16, 168]. The supplementation of potassium, magnesium, zinc, vitamins C, D, B6, and a decreased intake of sodium and selenium can "positively modulate blood pressure levels" [32]. The aforementioned study involving responses to a head up tilt test in vitamin B12 deficient people found a drop in systolic blood pressure 60 beats after the test [16]. This finding aligns with previous work suggesting that a dip in blood pressure when standing up from sitting or lying down is a symptom of vitamin B12 deficiency [168].
2.3.2 Sleep
Another area of research is the role of micronutrient status in sleep. Sleep duration is associated positively with iron, zinc, and magnesium and negatively with copper, potassium, vitamin A and vitamin B12 levels [74, 18]. Sleep quality increases with zinc, magnesium, and vitamin B9 status and is negatively associated with vitamin B12 status [31, 71, 18]. There are conflicting findings for iron. One study reports that iron status is not proven to be correlated with sleep quality [73], while another claims that supplementation had positive effects on sleep disorders [98].
2.3.3 General Symptoms
@@
Deficiencies can be classified as either clinical or subclinical based on their severity [167]. Most deficiencies result in symptoms of general fatigue, lethargy, irritability, muscle pain, weakness, and headaches. Clinical deficiencies often have more distinguishable symptoms, while subclinical deficiencies are limited to the above non-specific ones. Deficiencies of vitamin C, B vitamins, iron, magnesium, and zinc have been linked to fatigue more so than others [167, 9]. While energy and fatigue are more subjective and can can rely on subject-reporting, there are some established and validated assessment methods such as the Multidimensional Fatigue Inventory [158] and the SF-36 Vitality Scale [181]. A comprehensive reference for physiological symptoms associated with deficiency is lacking in literature, so we provide one in Tables LABEL:tab:phys-vitamins-water-soluble, LABEL:tab:phys-vitamins-fat-soluble, and LABEL:tab:phys-minerals within the Appendix.
3 Biofluid Analysis Methods for Micronutrient Status Assessment

### v2

@@ 2.1. Micronutrient Characteristics
Micronutrients are divided into three categories: water-soluble vitamins, fat-soluble vitamins, and minerals (Callahan et al., 2020). Water-soluble vitamins, such as vitamin C and the B vitamins, are absorbed directly into the bloodstream and are quickly excreted in urine. They are not stored for long periods in the body, so regular intake is necessary to prevent deficiencies, and there is less concern about toxicity from excess intake. On the other hand, fat-soluble vitamins like A, D, E, and K are absorbed into lymph vessels along with dietary fats. Fat-soluble vitamins are stored in larger quantities in fatty tissues and the liver, so deficiencies take longer to develop, and daily intake is less critical. However, due to their efficient storage and the lack of a rapid excretion mechanism, toxicity is more of a concern.
@@
Minerals can be categorized as major or trace minerals based on the daily requirement (Callahan et al., 2020). Major minerals, such as sodium, potassium, chloride, phosphorus, and magnesium, are required in amounts greater than 100 mg per day, while trace minerals like iron, copper, zinc, selenium, iodine, chromium, fluoride, and manganese are needed in amounts of 100 mg or less per day. Minerals are water-soluble and are absorbed directly into the bloodstream, sometimes with the help of transport proteins. It is important to note that minerals have an electric charge, and their function and storage can be influenced by various factors. For example, certain minerals carry a positive charge, such as sodium (Na+) and potassium (K+), while others are negatively charged, like chloride (Cl-). Their electric charge plays a crucial role in mineral homeostasis, influencing absorption efficiency, competition for transport and storage, bioavailability, and electrolyte balance. For a more in-depth discussion of these interactions, we refer the reader to (Solomons, 1986; Varvara and Vodnar, 2023). A comprehensive summary of the characteristics of micronutrients is presented in the Appendix in Tables LABEL:tab:char-vitamins-water-soluble, LABEL:tab:char-vitamins-fat-soluble, and LABEL:tab:char-minerals.
@@
Dietary guidelines for intake vary across organizations and countries, and are based on factors such as age and sex. These guidelines are not standardized due to the individualized nature of micronutrient metabolism, the diversity of micronutrients, their interactions, and the specific values at which they are needed. The US, for example, utilizes dietary reference intervals such as the Recommended Dietary Allowance (RDA) (National Institutes of Health, dateb). The RDA represents the average daily level of intake sufficient to meet the nutrient requirements of 97-98% of healthy individuals. The US National Institutes of Health (NIH) also defines terms to describe intake levels, such as an Upper Limit (UL), highlighting the potential for excess intake. In cases where there is insufficient evidence to establish these guidelines, there exists the more general term, Adequate Intake, which establishes a lower bound of nutritional intake necessary to meet a healthy nutritional state within a population (Institute of Medicine Subcommittee on Interpretation and Uses of Dietary Reference Intakes and Institute of Medicine (US) Standing Committee on the Scientific Evaluation of Dietary Reference Intakes(2000), US). For example, Adequate Intake for the nutrient biotin (vitamin B7) in human milk-fed infants is defined by the biotin content of human milk itself, because there is a lack of data available to scientifically determine an RDA for this population. Adequate Intake further emphasizes the uncertainty surrounding the appropriate levels of micronutrients in diet.
Micronutrient status assessment usually relies on the combination of the analysis of micronutrient biomarkers in biofluids as well as the physiological symptoms presented by an imbalance. Micronutrient biomarkers (predominantly measured by laboratory analyses of biofluids) are still debated, but can generally serve as a reliable optimal reference standard for assessment methods (Elmadfa and Meyer, 2014). An optimal reference standard is the clinically agreed-upon technique for determining a patient’s in-body status of a micronutrient (e.g. vitamin B9), and describes the target biomarker (e.g. folate), biofluid matrix to be analyzed (e.g. blood serum), and analysis method (e.g. LC-MS/MS) to use for accurate assessment. These standards are continuously debated by the clinical community, and are typically defined by current knowledge of “the chemistry, absorption, distribution in the body, and metabolism” of a nutrient (Elmadfa and Meyer, 2014). Because these dynamics can be highly individualized, validating a standard requires well-controlled studies that involve specific dietary interventions and the subsequent evaluation of the standard’s “specificity, sensitivity, and suitability for various population subgroups” (Elmadfa and Meyer, 2014).
Certain micronutrient imbalances are associated with observable physical or physiological symptoms, such as those in the skin, eyes, autonomic functioning, etc., which may present opportunities for developing noninvasive assessment methods tailored to specific nutrient deficiencies (Hummell and Cummings, 2022; Mauldin et al., 2021; Esper, 2015; DiBaise and Tarleton, 2019; Radler and Lister, 2013). These symptoms provide noninvasive insights into the bodily storage of micronutrients, internal processes, and can help identify appropriate biomarkers for further testing. Section 2.3 discusses micronutrient effects on physiological processes in more depth. It should be noted that some physical symptoms of micronutrient imbalances only arise when the imbalance becomes quite severe. This can happen slowly, in some cases over the course of several months (i.e. iron), and can result in irreversible symptoms (e.g. night blindness as a result of severe vitamin A deficiency; Tables LABEL:tab:_char-ws - LABEL:tab:_char-min). While physical symptoms still provide critical insights into the overall burden of micronutrient imbalance in a community, sole reliance on them for assessment can complicate preventative treatment. Quantitative, biofluid-based assessments for biomarkers are not immune to similar issues faced by physiological assessments (e.g. status of the biomarker plasma retinol decreases only after vitamin A stores in liver and eyes have nearly depleted), emphasizing the importance of considering both physiological and biofluid-based assessments in the practice of clinical nutrition.
2.2. Micronutrients in Biofluids
@@
Biofluid analysis for micronutrient status remains challenging since it is often unclear how other biofluids reflect the optimal reference standard matrix for that micronutrient. Some biofluids include blood, saliva, sweat, tears, urine. Besides blood, each of these can be collected and analyzed noninvasively. Blood and urine are clinically relevant for representing in-body micronutrient status, as will be demonstrated in Section 3.1. However, evidence is less clear for saliva, sweat, and tears (Höller et al., 2018). Although it will not be covered in depth, it is also worth mentioning human milk as a potential biofluid. Breast milk may have implications on the developmental outcomes of infants whose primary source of nutrition is human milk, but this is actively debated (Lockyer et al., 2021). As mentioned earlier, iodine excess in breastfeeding parents can also express itself in human milk, placing infants at an increased risk (Pike and Zlotkin, 2019).
In the clinical literature, the equivalence of saliva, sweat, and tears to blood and urine in micronutrient status assessment remains ambiguous. Some evidence for correlation of micronutrient levels in saliva with blood levels was found on a by-micronutrient basis (Höller et al., 2018). For example, serum and saliva levels of vitamin D were found to have a correlation of 0.56, measured using a total vitamin D (25-hydroxy vitamin D) kit with the electrochemiluminescence technique (Bahramian et al., 2018). Additionally, some correlation of iron levels in saliva and serum (a blood derivative) have been found. However, validity remains inconclusive as some sources have found a high positive correlation between salivary and serum levels (Canatan and Akdeniz, 2012; Haji-Sattari et al., 2019), while others report a high negative correlation (Flora et al., 2021; B et al., 2021; Gawaly et al., 2020). The techniques used to study this correlation involve enzyme-linked immunosorbent assays (ELISAs), laboratory assays, spectrophotometry, and chemiluminescence methods.
Additional research aims to quantify micronutrient status with saliva as well as tears (Ruiz-Valdepeñas Montiel et al., 2021; Sempionatto et al., 2019). The three types of human tears, basal, reflex, and emotional, differ in their chemical composition, which can influence which type is most appropriate to collect for specific analyses (Liang et al., 2023). Sempionatto et al. (2019) in particular argue that tears are a good biofluid for analysis since they are noninvasive, less complex than blood yet still contain a ”variety of biomarkers”, and they ”reflect concurrent blood levels” because of passive leakage of compounds from blood plasma. It is important to note, however, that neither of these works ((Ruiz-Valdepeñas Montiel et al., 2021; Sempionatto et al., 2019)) provide comparisons to the in-body status of their target micronutrients as measured by optimal reference standard, urine or blood-based assays. Additionally, it is important to consider the potential ethical implications of tear inducement and collection. One study used Schirmer strips, a common method for collecting tears in healthcare settings, and found that most participants considered the process acceptable (Quah et al., 2014). Specifically, 70% did not mind the procedure, and 74% preferred tear collection over venous blood sampling or other forms of biofluid collection such as urine.
@@
Sweat receives significant attention in emerging micronutrient detection methods, especially in wearables. The composition of collected sweat varies depending on whether it is produced actively (through exercise or heat exposure) or passively (through methods that encourage sweat production without physical activity, such as sweat patches or chemical stimulation) (Klous et al., 2020). Active sweating in particular usually produces higher concentrations of sodium, chloride, and metabolites. Eccrine sweat is a clear, odorless fluid secreted by eccrine sweat glands, which are essential for thermoregulation (Hodge et al., 2022). It is the primary sweat type used for micronutrient analysis, as apocrine sweat—the other major type—is thicker and predominantly composed of lipids and proteins, making it less suitable for this purpose (Heikenfeld et al., 2023). Therefore, when we discuss sweat, we refer to eccrine sweat, which is mostly sodium and chloride, with smaller amounts of micronutrients and metabolites (at the micro and nanomolar scale), similar to or smaller than their concentrations in blood plasma (Baker and Wolfe, 2020). Micronutrients found in sweat include potassium, calcium, magnesium, iron, copper, zinc, vitamin C, and vitamin B1. Out of the compounds in sweat, mostly sodium and chlorine ions are well-studied. Some research has explored water-soluble vitamins in sweat, such as vitamin C and vitamin B1, but there is no such attention on fat-soluble vitamins. Many confounding factors can impact sweat composition, such as the contamination of sweat by skin-derived substances, notably iron (Baker and Wolfe, 2020). This is combated by pre-rinsed skin, removal of initial sweat (concentrations stabilize after 20 to 30 minutes of sweating), and the analysis of cell-free sweat. The region and method of collection can also have an impact, as many micronutrient concentrations can vary two to four times depending on the region. Finally, sweat can be reabsorbed into the body. Skin temperature and the flow rate of sweat both impact the rate of this reabsorption.
@@
The clinical literature notes a general ”lack of association between dietary micronutrient intake and corresponding sweat micronutrient concentrations” (Baker and Wolfe, 2020). This lack of an association exists in comparison to blood as well. A review by Baker and Wolfe (2020) finds that there is no established correlation between sweat and blood composition, and there is ”little support for using sweat as a surrogate for blood”. Concentrations of minerals in sweat are much more varied than in plasma, likely because minerals bind to carrier proteins in blood. The review also reports little to no correlation between sweat and blood concentrations of vitamin C and iron status. This finding for iron is echoed in another paper that found both iron and calcium have no correlation between sweat and blood concentrations (Baker, 2019). However, it was reported in the same paper that iron concentrations in sweat have been observed to be lower in anemic patients and higher in patients undergoing iron therapy.
Finally, it is important to consider the potential impact of time lag variability in biomarker expression across different body fluids. This variability arises from nutrient kinetics and bioaccessibility, particularly in relation to absorption, distribution, metabolism, and excretion (ADME) within the human body (Mengucci et al., 2020; Gonzalez et al., 2023). However, studies described in this section do not explicitly account for these temporal effects, which are crucial for accurately correlating blood biomarkers with those present in other biofluids. At present, the temporal dynamics governing these relationships remain insufficiently understood, and require further investigation.
2.3. Micronutrient Effects on Physiological Processes
@@
Most of the physiological effects of micronutrients are related to the autonomic functions of the body. Autonomic functioning refers to bodily functions controlled by the autonomic nervous system, which regulates involuntary processes and rhythms such as breathing, heart rate, and digestion (Waxenbaum et al., 2023). The specific importance of micronutrients to the proper functioning of the nervous system has been documented by several papers (Georgieff, 2007; Beitzke et al., 2002; Chiu et al., 2021). For example, a study of vitamin B12 deficiency compared responses to a 60-degree passive head up tilt test between a control group, a vitamin B12 deficient group, and a group with diabetes mellitus (Beitzke et al., 2002). They found that the deficient group had comparable autonomic neuropathy to the diabetic group. Our exploration of observable impacts of deficiencies on autonomic functioning found three main affected areas relating to biofluid composition and physiological effects: general symptoms, cardiac function, and sleep.
2.3.1. General Symptoms
@@
Deficiencies can be classified as either clinical or subclinical based on their severity (Tardy et al., 2020). Most deficiencies result in symptoms of general fatigue, lethargy, irritability, muscle pain, weakness, and headaches. Clinical deficiencies often have more distinguishable symptoms, while subclinical deficiencies are limited to the above non-specific ones. Deficiencies of vitamin C, B vitamins, iron, magnesium, and zinc have been linked to fatigue more so than others (Tardy et al., 2020; Azzolino et al., 2020). While energy and fatigue are more subjective and can can rely on subject-reporting, there are some established and validated assessment methods such as the Multidimensional Fatigue Inventory (Smets et al., 1995) and the SF-36 Vitality Scale (Ware and Kosinski, 2001). A comprehensive reference for physiological symptoms associated with deficiency is lacking in literature, so we provide one in Tables LABEL:tab:phys-vitamins-water-soluble, LABEL:tab:phys-vitamins-fat-soluble, and LABEL:tab:phys-minerals within the Appendix.
2.3.2. Cardiac Function
One main aspect of autonomic functioning affected by nutrient imbalance is cardiac functioning. Several studies explore the interaction between heart rate variability (HRV) and micronutrient deficiencies. Components of HRV are associated with parasympathetic (PNS) and sympathetic nervous system (SNS) activity (Shaffer and Ginsberg, 2017). High-frequency (HF) bands reflect PNS activity and correspond to the respiratory cycle, while low-frequency (LF) bands reflect PNS, SNS, and baroreceptor activity. Vitamin B12 deficiency is one of the most documented in terms of impact to HRV, with evidence that it lowers HRV overall, impacting sympathetic indices the most (Aytemir et al., 2000; Sözen et al., 1998; Lopresti, 2020a; Beitzke et al., 2002). Supplementation of B12 was also demonstrated to return HRV indices to a comparably normal state (Aytemir et al., 2000). Deficiency of vitamin D was found to lower HRV as well (Lopresti, 2020a). Calcidiol (25(OH)D) levels, a form of vitamin D, was shown to be associated with the ratio of LF to HF HRV power (Mann et al., 2013). This metric is sometimes called sympathovagal balance, and is intended to be a measure of ’balance’ between SNS and PNS activity, but there has been debate over this interpretation (Shaffer and Ginsberg, 2017). Iron-deficiency anemia (IDA), an advanced form of iron deficiency, has more conflicting evidence of HRV impacts, with some studies finding no difference in HRV indices versus the control (Tuncer et al., 2009) while others were able to find a difference in the IDA group (Jibhkate and Lath, 2019; Yokusoglu et al., 2007).
Impacts of micronutrients on blood pressure have also been studied (Chiu et al., 2021; Beitzke et al., 2002; Toru et al., 1999). The supplementation of potassium, magnesium, zinc, vitamins C, D, B6, and a decreased intake of sodium and selenium can ”positively modulate blood pressure levels” (Chiu et al., 2021). The aforementioned study involving responses to a head up tilt test in vitamin B12 deficient people found a drop in systolic blood pressure 60 beats after the test (Beitzke et al., 2002). This finding aligns with previous work suggesting that a dip in blood pressure when standing up from sitting or lying down is a symptom of vitamin B12 deficiency (Toru et al., 1999).
As an aside, pulse-oximiters, smartwatches, and other health sensors or even smartphones can readily measure continuous cardiac function through photoplethysmography (PPG). Pulse rate (PR), HRV, and blood pressure can be derived from PPG (Mousavi et al., 2019; Liu et al., 2021).
2.3.3. Sleep
Another area of research is the role of micronutrient status in sleep. Sleep duration is associated positively with iron, zinc, and magnesium and negatively with copper, potassium, vitamin A and vitamin B12 levels (Ji et al., 2017; Beydoun et al., 2014). Sleep quality increases with zinc, magnesium, and vitamin B9 status and is negatively associated with vitamin B12 status (Cherasse and Urade, 2017; Jahrami et al., 2021; Beydoun et al., 2014). There are conflicting findings for iron. One study reports that iron status is not proven to be correlated with sleep quality (Ji, 2017), while another claims that supplementation had positive effects on sleep disorders (Leung et al., 2020). Sleep deprivation is also connected to micronutrition through its influence on hormones that regulate stress and the immune system. Sleep deprivation can decrease levels of cortisol while increasing ghrelin, a hormone linked to hunger (Kim et al., 2015b). This shift can lead to increased hunger, and a continuous lack of sleep was even found to have a positive correlation with obesity (which is interconnected with malnutrition (World Health Organization, 2024)). Studies also found that after temporary sleep deprivation there were decreased magnesium levels measured via red blood cell testing and reduced zinc levels in plasma tests (Lopresti, 2020b).
3. Biofluid Analysis Methods for Micronutrient Status Assessment